The point with position vectors (2,6) (1,2) (p,10) are collinear if the value of p is

Grade:12

1 Answers

Vikas TU
7 years ago
If the three points are ccollinear the the slope of the line passing these three points would be same.
Therfore,
(10-2)/(p-1) = (2-6)/(1-2)
=> 8/(p-1) = 4
=> p-1 = 2
p = 3